# LogicTree: Structured Proof Exploration for Coherent and Rigorous Logical Reasoning with Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have achieved remarkable multi-step reasoning capabilities across various domains. However, LLMs still face distinct challenges in complex logical reasoning, as (1) proof-finding requires systematic exploration and the maintenance of logical coherence and (2) searching the right combination of premises at each reasoning step is inherently challenging in tasks with large premise space. To address this, we propose LogicTree, an inference-time modular framework employing algorithm-guided search to automate structured proof exploration and ensure logical coherence. Advancing beyond tree-of-thought (ToT), we incorporate caching mechanism into LogicTree to enable effective utilization of historical knowledge, preventing reasoning stagnation and minimizing redundancy. Furthermore, we address the combinatorial complexity of premise search by decomposing it into a linear process. The refined premise selection restricts subsequent inference to at most one derivation per step, enhancing reasoning granularity and enforcing strict step-by-step reasoning. Additionally, we introduce two LLM-free heuristics for premise prioritization, enabling strategic proof search. Experimental results on five datasets demonstrate that LogicTree optimally scales inference-time computation to achieve higher proof accuracy, surpassing chain-of-thought (CoT) and ToT with average gains of 23.6% and 12.5%, respectively, on GPT-4o. Moreover, within LogicTree, GPT-4o outperforms o3-mini by 7.6% on average.
@@ 1 Introduction
To tackle these challenges, some studies use an iterative framework to build longer reasoning chains for solving complex problems  (Creswell et al., 2023). Within the framework, they adopt a modular approach to decompose individual reasoning steps  (Khot et al., 2023), separating premise selection from inference and assigning each to specialized LLM modules for improved accuracy  (Xu et al., 2024b; Zhang et al., 2024; Sun et al., 2024). Further research integrates LLM modules into tree structures, enabling systematic proof exploration  (Kazemi et al., 2023; Yao et al., 2023; Wang et al., 2025). Although these methods have achieved notable advancements, there are still limitations:
@@
(1) Difficulty in maintaining logical coherence and less effective utilization of derived knowledge obstruct progressive proof construction. In some iterative approaches  (Creswell et al., 2023; Zhang et al., 2024), reasoning steps are not required to build directly on prior derivations, which may disrupt logical coherence, hinder deep reasoning and cause redundancy  (Wang et al., 2025). While tree-based method  (Yao et al., 2023) mitigates this issue, they lack mechanisms to utilize derived facts across branches, potentially leading to reasoning stagnation  (Sun et al., 2024).
@@
To address these limitations, we propose LogicTree, a novel inference-time modular framework for structured proof exploration. The overview of our framework is shown in Figure 1. LogicTree includes four LLM-based modules: Forward Selection, Backward Selection, Derivation and Verification, which are embedded in tree structure. Additionally, we incorporate Fact Repository and Derivation HashMap into LogicTree, functioning as caching mechanism. Fact Repository is initialized with given facts and dynamically stores derived facts. It enables branches to access the given facts and the derivations from earlier branches, facilitating cross-branch information flow and effective utilization of historical knowledge throughout proof exploration. Derivation HashMap records derived facts along with their derivation paths for a traceable reasoning process. We employ depth-first search (DFS) to orchestrate LLM modules and cache components, automating systematic proof search while ensuring logical coherence.
@@ 2.2 Inference-time Scaling for Reasoning
Modular inference without LLM parameter updates. Modular approach decomposes complex reasoning tasks into simpler sub-tasks, each assigned to specialized LLM modules implemented through few-shot prompting  (Khot et al., 2023). In logical reasoning, it involves two key modules that operate iteratively: premise selection and inference  (Creswell et al., 2023). Extending from this foundation, Cumulative Reasoning  (Zhang et al., 2024) integrates LLM verifier to validate reasoning steps. DetermLR  (Sun et al., 2024) employs LLM scorer to prioritize relevant premises. SymbCoT  (Xu et al., 2024b) and Aristotle  (Xu et al., 2024a) introduce LLM translator to convert natural language input into symbolic representations. Further research embeds LLM modules into topological structures, enabling deliberate problem solving  (Yao et al., 2023; Besta et al., 2024).
Current LLM-based methods for logical reasoning still struggle with structured exploration, maintaining logical coherence, and ensuring reasoning rigor in complex, multi-step tasks. To address these challenges, we propose a novel inference-time modular approach that enables systematic and extensive proof exploration and enhances reasoning rigor.
@@ 3.1 Task Definition
Logical reasoning aims to determine the truth value (true, false, or unknown) of a hypothesis $\mathcal{H}$ based on a set of premises consisting of facts $\mathcal{F}$ and rules $\mathcal{R}$  (Dowden, 2020). An example is shown in Figure 5. Formally, =$\mathcal{F}=\{f_{i}\mid i=1,2,\dots,N_{\mathcal{F}}\}$, where each $f_{i}$ represents a definitive statement within the reasoning system. =$\mathcal{R}=\{r_{i}\mid i=1,2,\dots,N_{\mathcal{R}}\}$, where each $r_{i}$ represents a conditional statement that defines a logical relationship between facts and inferred conclusions. The reasoning process applies standard logical operators, including: Negation ($\neg$), Conjunction ($\land$), Disjunction ($\lor$), Implication ($\Rightarrow$), Equivalence ($\Leftrightarrow$). We define the set of intermediate derived facts as $\mathcal{D}=\{d_{i}\mid i=1,2,\dots,N_{\mathcal{D}}\}$.
@@ 3.2 Components of LogicTree Framework
As shown in Figure 1, the LogicTree framework includes (1) two caches: Fact Repository and Derivation HashMap; (2) four LLM-based modules: Forward Selection, Derivation, Backward Selection, and Verification, each implemented by few-shot prompting. The specific prompts for each module, along with example inputs and outputs, are provided in Appendix H.
@@
Derivation Module. Along each branch, this module performs a strict one-step derivation using the current leaf rule and its parent fact. A successful derivation occurs if the fact fully satisfies the rule’s condition. The newly derived fact is added as a child node to the tree only if it is not already present in Fact Repository (to avoid redundancy). If the derivation fails, it results from one of the two reasons: (1) the fact does not satisfy the rule’s condition at all, i.e., the rule was incorrectly selected by Forward Selection Module; (2) the fact partially satisfies the rule’s conditions, with some required fact(s) still missing. In the first case, the branch is marked as a dead end. In the second case, where conjunctive reasoning (e.g., $f_{1}\land f_{2}\land r_{1}\Rightarrow\ d_{1}$) is required, the branch is marked as a pseudo dead-end, where the missing fact(s) may still be retrievable.
@@
Verification Module. This module evaluates the derived fact $d_{i}$ against the hypothesis $\mathcal{H}$ to determine if the proof is complete. If the derived fact is equivalent to or directly contradicts the hypothesis, the proof is concluded; otherwise, proof exploration continues.
@@ 3.3 LLM-free Premise Prioritization
Fact (root) ranking for global ordering of tree search. In LogicTree framework, each given fact $f_{i}$ serves as the root of a tree, and trees are explored sequentially until the proof is found. As shown in Figure 1a, we first apply a semantic alignment step to prioritize facts that have the same subject with the hypothesis $\mathcal{H}$ as tree roots. To further rank facts, we define cumulative connectivity between a fact $f_{i}$ and the rule set $\mathcal{R}$, which is the sum of semantic overlap between $f_{i}$ and each rule $r_{i}\in\mathcal{R}$. It approximates how many reasoning branches the root fact can initiate through its relevant rules. Facts with zero connectivity are discarded, as they cannot contribute to any derivation. Those with higher connectivity are prioritized for opening more reasoning paths and higher likelihood of proof discovery in earlier-explored trees. We conduct subject alignment and compute semantic overlap using spaCy’s efficient dependency parsing.
Rule ranking at local level for early stopping. Rule ranking sorts the rules selected by Forward Selection Module based on each rule’s semantic overlap with hypothesis $\mathcal{H}$. This prioritization directs Derivation Module to first apply the rule $r_{i}$ whose derivation is most likely to verify hypothesis $\mathcal{H}$, facilitating early stopping. For example, to verify $\mathcal{H}$: "Kevin is uncomfortable.", a rule $r_{i}$ such as: "If …, the person is uncomfortable." is prioritized.
@@ 3.4 LogicTree Algorithm
After each inference, Verification evaluates the result to determine whether it concludes the proof, enabling early stopping (Algorithm 3). If the result indicates an underivable or redundant outcome (i.e., the branch reaches a dead end), DFS backtracks to explore the next branch. Otherwise, the derived fact $d_{i}$ is appended to the tree for further expansion.
The next iteration begins from derived fact $d_{i}$, with LLM modules reset before the next call. By building each step upon prior derivations, our framework maintains logical coherence. Instead of restricting the tree’s depth or breadth, we set a step limit. Once a tree is fully explored, the algorithm proceeds to the next tree, using the next ranked fact as the root. If all trees are explored or the step limit is reached without verifying the hypothesis $\mathcal{H}$, the algorithm returns Unknown.
@@ 4.1 Experimental Setup
Datasets. We evaluate our framework on five multi-step logical reasoning datasets: RobustLR  (Sanyal et al., 2022), PrOntoQA-OOD  (Saparov et al., 2023), ProofWriter  (Tafjord et al., 2020), ParaRules  (Clark et al., 2020), LogicNLI  (Tian et al., 2021). For all examples in our experiments, hypothesis concludes as True, False, or Unknown. More details are provided in Appendix C.1.
@@
  • Modular approaches: SI (Selection-Inference)  (Creswell et al., 2023) adpots selection and inference modules for iterative reasoning. CR (Cumulative Reasoning)  (Zhang et al., 2024) introduces a cumulative process of generating new propositions to reach the answer. ToT (Tree-of-Thought)  (Yao et al., 2023) leverages tree-search algorithm for deliberate reasoning. LAMBADA  (Kazemi et al., 2023) develops a backward chaining approach for automated reasoning.
@@
Models. Our proposed framework places no restrictions on the choice of LLMs. Here, we separately employ GPT-4o-mini, GPT-4o,  (Achiam et al., 2023) and Llama-3.3 70B  (Dubey et al., 2024) within our framework. For a fair comparison, we reproduce CoT and other modular approaches using the same models. Further details on models are provided in Appendix C.2.
Evaluate reasoning accuracy. In logical reasoning, correct label prediction (True, False, or Unknown) does not necessarily indicate correct reasoning, as models may arrive at the correct conclusion through hallucinated premises or spurious correlations  (Kazemi et al., 2023; Liu et al., 2023). Similar to  Saparov and He (2023), we use proof accuracy for rigorous evaluation. We manually verify each example by focusing on the reasoning chain that verifies the hypothesis within the full reasoning trace. A proof is considered correct if every step in this chain is valid, while the validity of unrelated reasoning attempts is disregarded.
@@ 4.2 Main Results
As shown in Table 3.4, our proposed LogicTree consistently outperforms CoT and other modular approaches across all five datasets. Specifically, our method surpasses CoT significantly, with average performance gains of 24.5%, 23.6%, and 19.6% on GPT-4o-mini, GPT-4o, and Llama-3.3 70B, respectively. Compared to ToT, the strongest among other modular methods, our framework achieves average improvements of 9.2%, 12.5%, and 9.8% on the same models. On ParaRules, PrOntoQA-OOD, ProofWriter, and RobustLR datasets, our framework achieves near-perfect proof accuracy with GPT-4o, highlighting its strength in logical reasoning. This strength generalizes across different levels of task difficulty, as shown in Figure 7. Furthermore, LogicTree demonstrates superior performance over RL-trained reasoning models, o1-mini and o3-mini, as presented in Figure 2. With Llama-3.3 70B, our method yields 8.7% and 5.2% higher accuracy on average; with GPT-4o, it achieves average gains of 11.1% and 7.6%, respectively.
@@ 5.1 Factors Impacting Proof Accuracy
Our framework adopts forward reasoning to leverage its aforementioned advantage, while decomposing premise selection to address its search complexity (Table 3), effectively improving both selection accuracy and inference accuracy, as shown in Figure 3b. Another key reason our framework outperforms ToT is that ToT lacks a mechanism to utilize derived facts from earlier branches, which may lead to reasoning stagnation (the scenario in Figure 1b-2). Our framework addresses this issue through incorporating Fact Repository (§ 3.2). The impact of this component is evaluated in Table 5.
@@ 5.2 Scaling of Reasoning Length
Proof accuracy vs. reasoning steps. To assess the impact of long reasoning on solving complex logical tasks and compare its effectiveness across different approaches, we measure the average number of reasoning steps for each approach across five datasets (details in Appendix E Table H). The corresponding proof accuracy and average reasoning steps for each approach are presented in Figure 4. Insufficient reasoning of CoT in complex tasks limits its performance. SI and CR, as analyzed in § 5.1, suffer from high proportion of redundant and failed inferences, which undermine the effectiveness of long reasoning. LAMBADA (backward reasoning) demonstrates more reasoning steps and lower proof accuracy compared to forward reasoning (ToT and LogicTree). Additional analysis comparing forward and backward reasoning is provided in Appendix F Table 4. Compared to ToT, our framework requires more reasoning steps for two main reasons: (1) Our framework decomposes the combinatorial premise search, leading to more steps; (2) In ToT, multiple derivations can occur within a single step (Figure 6d), whereas our framework restricts each step to at most one derivation. The enhanced reasoning granularity ensures strict step-by-step reasoning, optimally increasing reasoning length and leveraging additional inference-time computation to achieve higher proof accuracy, as shown in the shaded areas of Figure 4 and Figure 8.

### v2

@@ Abstract
Large language models (LLMs) have achieved remarkable multi-step reasoning capabilities across various domains. However, LLMs still face distinct challenges in complex logical reasoning, as (1) proof-finding requires systematic exploration and the maintenance of logical coherence and (2) searching the right combination of premises at each reasoning step is inherently challenging in tasks with large premise space. To address this, we propose LogicTree, an inference-time modular framework employing algorithm-guided search to automate structured proof exploration and ensure logical coherence. Advancing beyond tree-of-thought (ToT), we incorporate caching mechanism into LogicTree to enable effective utilization of historical knowledge, preventing reasoning stagnation and minimizing redundancy. Furthermore, we address the combinatorial complexity of premise search by decomposing it into a linear process. The refined premise selection restricts subsequent inference to at most one derivation per step, enhancing reasoning granularity and enforcing strict step-by-step reasoning. Additionally, we introduce two LLM-free heuristics for premise prioritization, enabling strategic proof search. Experimental results on five datasets demonstrate that LogicTree optimally scales inference-time computation to achieve higher proof accuracy, surpassing chain-of-thought (CoT) and ToT with average gains of 23.6% and 12.5%, respectively, on GPT-4o. Moreover, within LogicTree, GPT-4o outperforms o3-mini by 7.6% on average. [^1] Our code is available at https://github.com/kang-ml/LogicTree.
@@ 1 Introduction
To tackle these challenges, some studies use an iterative framework to build longer reasoning chains for solving complex problems  (Creswell et al., 2023). Within the framework, they adopt a modular approach to decompose individual reasoning steps  (Khot et al., 2023), separating premise selection from inference and assigning each to specialized LLM modules for improved accuracy  (Xu et al., 2024b; Zhang et al., 2025; Sun et al., 2024). Further research integrates LLM modules into tree structures, enabling systematic proof exploration  (Kazemi et al., 2023; Yao et al., 2023; Wang et al., 2025). Although these methods have achieved notable advancements, there are still limitations:
@@
(1) Difficulties in maintaining logical coherence and in effectively utilizing derived knowledge obstruct progressive proof construction. In some iterative approaches  (Creswell et al., 2023; Zhang et al., 2025), reasoning steps are not required to build directly on prior derivations, which may disrupt logical coherence, hinder deep reasoning and cause redundancy  (Wang et al., 2025). While tree-based method  (Yao et al., 2023) mitigates this issue, it lacks mechanisms to share derived knowledge across branches, potentially leading to reasoning stagnation  (Sun et al., 2024).
@@
To address these limitations, we propose LogicTree, a novel inference-time modular framework for structured proof exploration. The overview of our framework is shown in Figure 1. LogicTree includes four LLM-based modules: Forward Selection, Backward Selection, Derivation and Verification, which are embedded in tree structure. Additionally, we incorporate Fact Repository and Derivation HashMap into LogicTree as cache components. Fact Repository is initialized with given facts and dynamically stores derived facts. It enables branches to access the given facts and the derivations from earlier branches, facilitating cross-branch information flow and effective utilization of historical knowledge throughout proof exploration. Derivation HashMap records derived facts along with their derivation paths for a traceable reasoning process. We employ depth-first search (DFS) to orchestrate LLM modules and cache components, automating systematic proof search while ensuring logical coherence.
@@ 2.2 Inference-time Scaling for Reasoning
Modular inference without LLM parameter updates. Modular approach decomposes complex reasoning tasks into simpler sub-tasks, each assigned to specialized LLM modules implemented through few-shot prompting  (Khot et al., 2023). In logical reasoning, it involves two key modules that operate iteratively: premise selection and inference  (Creswell et al., 2023). Extending from this foundation, Cumulative Reasoning  (Zhang et al., 2025) integrates LLM verifier to validate reasoning steps. DetermLR  (Sun et al., 2024) employs LLM scorer to prioritize relevant premises. SymbCoT  (Xu et al., 2024b) and Aristotle  (Xu et al., 2024a) introduce LLM translator to convert natural language input into symbolic representations. Further research embeds LLM modules into topological structures, enabling deliberate problem solving  (Yao et al., 2023; Besta et al., 2024).
Current LLM-based methods for logical reasoning still struggle to perform structured exploration while ensuring logical coherence and rigor in complex, multi-step reasoning tasks. To address these challenges, we propose a novel inference-time modular approach that enables systematic and extensive proof exploration and enhances reasoning rigor.
@@ 3.1 Task Definition
Logical reasoning aims to determine the truth value (true, false, or unknown) of a hypothesis $\mathcal{H}$ based on a set of premises consisting of facts $\mathcal{F}$ and rules $\mathcal{R}$  (Dowden, 2020). An example is shown in Figure 5. Formally, $\mathcal{F}=\{f_{i}\mid i=1,2,\dots,N_{\mathcal{F}}\}$, where each $f_{i}$ represents a definitive statement within the reasoning system. $\mathcal{R}=\{r_{i}\mid i=1,2,\dots,N_{\mathcal{R}}\}$, where each $r_{i}$ represents a conditional statement that defines a logical relationship between facts and inferred conclusions. The reasoning process applies standard logical operators, including: Negation ($\neg$), Conjunction ($\land$), Disjunction ($\lor$), Implication ($\Rightarrow$), Equivalence ($\Leftrightarrow$). We define the set of intermediate derived facts as $\mathcal{D}=\{d_{i}\mid i=1,2,\dots,N_{\mathcal{D}}\}$.
@@ 3.2 Components of LogicTree Framework
As shown in Figure 1, the LogicTree framework includes (1) two caches: Fact Repository and Derivation HashMap; (2) four LLM-based modules: Forward Selection, Derivation, Backward Selection, and Verification, each implemented by few-shot prompting. The specific prompts for each module, along with example inputs and outputs, are provided in Appendix I.
@@
Derivation Module. Along each branch, this module performs a strict one-step derivation using the current leaf rule and its parent fact. A successful derivation occurs if the fact fully satisfies the rule’s condition. If the derivation fails, it results from one of the two reasons: (1) the fact does not satisfy the rule’s condition at all, i.e., the rule was incorrectly selected by Forward Selection Module; (2) the fact partially satisfies the rule’s conditions, with some required fact(s) still missing. In the first case, the branch is marked as a dead end. In the second case, where conjunctive reasoning (e.g., $f_{1}\land f_{2}\land r_{1}\Rightarrow\ d_{1}$) is required, the branch is marked as a pseudo dead-end, where the missing fact(s) may still be retrievable.
@@
Verification Module. After each successful derivation, this module evaluates the derived fact $d_{i}$ against the hypothesis $\mathcal{H}$ to determine if the proof is complete. If the derived fact is equivalent to or directly contradicts the hypothesis, the proof is concluded; otherwise, proof exploration continues.
@@ 3.3 LLM-free Premise Prioritization
Fact (root) ranking for global ordering of tree search. In LogicTree framework, each given fact $f_{i}$ serves as the root of a tree, and trees are explored sequentially until the proof is found. As shown in Figure 1a, we first apply a semantic alignment step to prioritize facts that have the same subject with the hypothesis $\mathcal{H}$ as tree roots. To further rank facts, we define cumulative connectivity between a fact $f_{i}$ and the rule set $\mathcal{R}$, which is the sum of semantic overlap between $f_{i}$ and each rule $r_{i}\in\mathcal{R}$. It approximates how many reasoning branches the root fact can initiate through its relevant rules. Facts with zero connectivity are discarded, as they cannot contribute to any derivation. Facts with higher connectivity are prioritized for opening more reasoning paths and higher likelihood of proof discovery in earlier-explored trees. We conduct subject alignment and compute semantic overlap using spaCy’s efficient dependency parsing.
Rule ranking at local level for early stopping.  After each Forward Selection, the selected rules are ranked based on each rule’s semantic overlap with hypothesis $\mathcal{H}$. This prioritization directs Derivation Module to first apply the rule $r_{i}$ whose derivation is most likely to verify hypothesis $\mathcal{H}$, facilitating early stopping. For example, to verify $\mathcal{H}$: "Kevin is uncomfortable.", a rule $r_{i}$ such as: "If …, the person is uncomfortable." would be prioritized.
@@ 3.4 LogicTree Algorithm
After each inference, Verification evaluates the result to determine whether it concludes the proof, enabling early stopping (Algorithm 3). If the result indicates an underivable or redundant (i.e., already in Fact Repository) outcome, DFS backtracks to explore the next branch. Otherwise, the derived fact $d_{i}$ is appended to the tree for further expansion.
The next iteration begins from derived fact $d_{i}$, with LLM modules reset before the next call. By building each step upon prior derivations, our framework maintains logical coherence. Once a tree is fully explored, the algorithm proceeds to the next tree, using the next ranked fact as the root. To avoid excessively long reasoning, we set an LLM query limit on the reasoning process. If all trees are explored or the query limit is reached (whichever occurs first) without verifying the hypothesis $\mathcal{H}$, the algorithm terminates and returns Unknown.
@@ 4.1 Experimental Setup
Datasets. We evaluate our framework on five multi-step logical reasoning datasets: RobustLR  (Sanyal et al., 2022), PrOntoQA-OOD  (Saparov et al., 2023), ProofWriter  (Tafjord et al., 2020), ParaRules  (Clark et al., 2020), LogicNLI  (Tian et al., 2021). For all examples in our experiments, hypothesis concludes as True, False, or Unknown. More details on logical reasoning datasets are provided in Appendix C.1. To further assess the broader adaptability of our framework, we extend it to mathematical reasoning datasets, as demonstrated in Appendix H.
@@
  • Modular approaches: SI (Selection-Inference)  (Creswell et al., 2023) adpots selection and inference modules for iterative reasoning. CR (Cumulative Reasoning)  (Zhang et al., 2025) introduces a cumulative process of generating new propositions to reach the answer. ToT (Tree-of-Thought)  (Yao et al., 2023) leverages tree-search algorithm for deliberate reasoning. LAMBADA  (Kazemi et al., 2023) develops a backward chaining approach for automated reasoning.
@@
Models. Our framework places no restrictions on the choice of LLMs. Here, we separately employ GPT-4o-mini, GPT-4o (Achiam et al., 2023), and Llama-3.3 70B  (Dubey et al., 2024) within our framework. We reproduce CoT and other modular approaches using the same models for comparison. Further details on models are in Appendix C.2.
Evaluate reasoning accuracy. In logical reasoning, correct label prediction (True, False, or Unknown) does not necessarily indicate correct reasoning, as models may arrive at the correct conclusion through hallucinated premises or spurious correlations  (Kazemi et al., 2023; Liu et al., 2023). Similar to  Saparov and He (2023), we use proof accuracy for rigorous evaluation. We manually verify each example by focusing on the reasoning chain that verifies the hypothesis within the entire reasoning trace. A proof is considered correct if every step in this chain is valid, while the validity of other reasoning paths is disregarded.
@@ 4.2 Main Results
As shown in Table 1, our proposed LogicTree consistently outperforms CoT and other modular approaches across all five datasets. Specifically, our method surpasses CoT significantly, with average performance gains of 24.5%, 23.6%, and 19.6% on GPT-4o-mini, GPT-4o, and Llama-3.3 70B, respectively. Compared to ToT, the strongest among other modular methods, our framework achieves average improvements of 9.2%, 12.5%, and 9.8% on the same models. On ParaRules, PrOntoQA-OOD, ProofWriter, and RobustLR datasets, our framework achieves near-perfect proof accuracy with GPT-4o, highlighting its strength in logical reasoning. This strength generalizes across different levels of task difficulty, as shown in Figure 7. Furthermore, when applied within LogicTree, Llama-3.3 70B and GPT-4o outperform RL-trained reasoning models, o1-mini and o3-mini, as shown in Figure 2. With Llama-3.3 70B, our method yields 8.7% and 5.2% higher accuracy on average; with GPT-4o, the average gains are 11.1% and 7.6%, respectively.
@@ 5.1 Factors Impacting Proof Accuracy
Our framework adopts forward reasoning to leverage its aforementioned advantage, while decomposing premise selection to address its search complexity (Table 3), effectively improving both selection accuracy and inference accuracy, as shown in Figure 3b. Another key reason our framework outperforms ToT is that ToT lacks a mechanism to leverage derived facts from earlier branches, which may lead to reasoning stagnation (the scenario in Figure 1b-2). Our framework addresses this issue through incorporating Fact Repository (§ 3.2). The impact of this component is evaluated in Table 5.
@@ 5.2 Scaling of Reasoning Length
Proof accuracy vs. reasoning length. To assess the impact of long reasoning on solving complex logical tasks and compare its effectiveness across different approaches, we measure the average number of reasoning steps for each approach across five datasets, with details in Appendix E. The corresponding proof accuracy and average reasoning steps for each approach are presented in Figure 4. Insufficient reasoning of CoT in complex tasks limits its performance. SI and CR, as analyzed in § 5.1, suffer from high proportion of redundant and failed inferences, which undermine the effectiveness of long reasoning. LAMBADA (backward reasoning) demonstrates more reasoning steps and lower proof accuracy compared to forward reasoning (ToT and LogicTree). Additional analysis comparing forward and backward reasoning is provided in Appendix F Table 4. Compared to ToT, our framework requires more reasoning steps for two main reasons: (1) Our framework decomposes the combinatorial premise search, leading to more steps; (2) In ToT, multiple derivations can occur within a single step (Figure 6d), whereas our framework restricts each step to at most one derivation. The enhanced reasoning granularity ensures strict step-by-step reasoning, optimally increasing reasoning length and leveraging additional inference-time computation to achieve higher proof accuracy, as shown in the shaded areas of Figure 4 and Figure 8.
We additionally report the average token cost and inference time for all methods in Table 7, complementing the reasoning-step counts. Figure 9 further illustrates the trade-off between accuracy and inference cost by showing proof accuracy under varying token budgets. The results indicate that (1) LogicTree consistently outperforms the baseline approaches at every budget level, and (2) LogicTree achieves larger performance gains as the token budget increases, demonstrating more effective scaling with additional inference computation.